# A model-agnostic gravitational-wave background characterization algorithm

### v1

@@ IV Flexible, Transdimensional Models
We propose an interpolation model to recover a GWB signal by placing a discrete number of points in an $n$-dimensional space and interpolating between these points. This requires a choice of interpolation scheme, which we describe in Section IV.1. The general interpolation model naturally lends itself to using a reversible jump Markov chain Monte Carlo (RJMCMC) algorithm [41] that allows the number of knots and thus the number of parameters in our model to vary. So, we must define how these interpolating points, or knots, move for each proposal in the RJMCMC. These knot movements are our proposals and are defined in Section IV.2. This is similar in spirit to the BayesLine and BayesWave algorithms [86, 87, 88].
@@ IV.1 Interpolation models
Instead of using a specific functional form for the GWB, such as a PL, we instead estimate the GWB at a set of chosen frequencies, which we call “knots,” and interpolate from those knots to the frequency bins at which measurements are taken. The parameters of the model are a set of $n$ amplitudes, $\mathbf{a}_{n}$, located at a set of knots $\mathbf{x}_{n}$, in frequency space. We use $x$ here to denote the location of a knot, to distinguish them from the frequency bins at which $\hat{C}(f)$ is calculated.
@@
with $f\in[x_{j},x_{j+1}]$. We refer to this model as the “linear interpolation” model or, if we perform the interpolation in log-space (i.e. $x\rightarrow\ln x$ and $a\rightarrow\ln a$ etc.) the “piece-wise power law” (PPL) model because linear interpolation in log-space corresponds to constructing the model from a set of PPLs. By working in log-space we can ensure that the GWB is positive (a defining characteristic of $\Omega_{\textrm{GW}}$), and that in the case of $n=1$ or $n=2$ knots the model simplifies to a constant or PL model. Moreover, for $n=2$, placing a log-uniform prior on $\mathbf{a}$ means that when we recover a PL, the prior we use is naturally the preferred prior advocated for in [48] and used in [48, 29].
@@
where the $c_{n,j}$’s are calculated using standard algorithms that solve for the $c_{n,j}$ subject to constraints that impose continuity and smoothness [89].
An additional interpolation method that we employ is the “Akima spline” model. Akima splines are non-smoothing splines that avoid the rapid oscillations between knots that are common in other spline formulations when the data are uninformative or the second derivative of the underlying function varies rapidly. An Akima spline model takes the same form as the cubic spline, but the coefficients, $c_{n,j}$, are informed by derivatives of neighboring knots instead of an optimization [90]. The result of an Akima spline interpolation is a smoother interpolation that is less sensitive to outliers in the data.
@@ IV.2 Reversible Jump Markov chain Monte Carlo
We start with an underlying grid of $N$ knots that are available to interpolate between. At any given time, $n$ of them will be used in the interpolation, or “activated”, and $N-n$ knots will be disregarded in the interpolation, or “deactivated”. The locations of the available underlying knots are labeled $x_{\alpha}$, where $\alpha=0,\ldots,N-1$. So, the initial location of $x_{\alpha}$ is given by:
@@
The parameters of our model are the list of amplitudes $\mathbf{a}_{n}$ corresponding to the amplitude of the model at the knot location $\mathbf{x}_{n}$ of each of the activated knots. We also allow the knot locations to vary, but maintain ordering on the underlying grid. That is, $x_{j}$ can vary within the range $f_{\textrm{min}}+2j\Delta x<x_{j}<f_{\textrm{min}}+2(j+1)\Delta x$ where $\Delta x=(f_{\textrm{max}}-f_{\textrm{min}})/(2N)$. So at any given time we have $2n$ parameters in our model. The prior on $\mathbf{x}_{n}$ and $\mathbf{a}_{n}$ throughout this paper is uniform within their allowed regions, and disjoint between knots  [^2] This simply implies two knots cannot be superposed.:
$p(\mathbf{a}_{n},\mathbf{x}_{n})=(2\Delta x)^{-n}(a_{\mathrm{max}}-a_{\mathrm{min}})^{-n}.$ (10)
We discuss the effect of these priors and how they change as the number of knots changes in App A.
@@
    Birth proposal: We propose to “turn on” one knot that is currently deactivated, choosing from the list of deactivated knots at random. This increases the number of parameters in the model and the overall prior volume. The location and amplitude of the proposed knot is either drawn from the prior, or the proposed $x_{j}^{new}$ is drawn from the prior, while $a_{j}$ is drawn from a Gaussian with a mean given by $\mathcal{G}(x_{j}^{new})$. In this case, the location and amplitudes of the knots for the interpolation, $x_{j}$ and $c_{n,j}$ in Eq. 6 or Eq. 8, are from the current sample. The standard deviation of the Gaussian is selected by the user when the sampler is instantiated (the default is $\sigma=1$).
@@
    Death proposal: We propose to “turn off” a knot that is currently active. This decreases the number of parameters in the model, and the overall prior volume.
@@
    Shift knot horizontally: we choose an active knot at random, and propose to shift its location, $x_{j}$ within its allowable region.
@@
We discuss the details of the algorithm in App. A, including the basics of RJMCMC, and more details about the implementation of the jump proposals.
Figure 1: A visualization of the RJMCMC implemented in this work. In both panels, we show the injected signal (black dashed) and the injected signal plus Gaussian white noise (blue points). The error bars on the blue points correspond to one standard deviation of the Gaussian white noise. In the top panel, the gray boxes correspond to the uniform priors that the interpolating knots (yellow stars) are allows to move within (see Eq. (10)). The knots are interpolating using cubic splines (red lines). The bottom panel shows the death proposal (see Sec. IV.2) where we have turned off one knot (gray star) and interpolated the remaining knots to fit the signal. This schematic highlights the transdimensional nature of the RJMCMC that we implement to fit injected GW signals with various detector sensitivities.
@@ V Detecting and Characterizing $\Omega_{\rm GW}(f)$
We choose three example injections to demonstrate our transdimensional spline interpolation model. These injections include a BBH-induced signal (Section V.1), a subtraction signal (Section V.2), and a signal based on a first-order phase transition model (Section V.3). For each example, after motivating the astrophysical relevance of each injection, we present the results of our injection and recovery. We perform the injection and recovery with both LIGO A+ and projected Cosmic Explorer (CE) sensitivity curves. For the CE sensitivity curve, we assume 40 km detector arms and the same location and orientation as the current LIGO interferometers at Hanford, WA, and Livingston, LA [42, 43].
For each injection, we simulate $\hat{C}(f)$ directly, by drawing $\hat{C}(f)\sim\mathcal{N}[\Omega_{\mathcal{M}}(f),\sigma(f)]$, where $\Omega_{\mathcal{M}}(f)$ is the injected model, and $\sigma(f)$ is the detector noise calculated using the specified noise curves and observation times. For each recovery, we perform $500,000$ Markov chain Monte Carlo steps, and discard the first $300,000$ as burn-in. We perform all interpolation in log-space, i.e. we take $x\rightarrow\ln x$, $a\rightarrow\ln a$, etc. when implementing the procedures outlined in Secs. IV.1 and IV.2.
@@ V.1 Full BBH injection
We use a BBH $\Omega_{\rm GW}$ as our first example of injection and recover it using this transdimensional spline interpolation approach. The contribution to the GWB from all BBHs can be written as
$\Omega_{\textrm{GW}}=\frac{f}{\rho_{c}}\int_{0}^{z_{\mathrm{max}}}\mathrm{d}z\;\frac{\mathcal{R}(z)\left<\frac{\mathrm{d}E}{\mathrm{d}f_{s}}\left.\right|_{f(1+z)}\right>}{(1+z)H(z)}\,,$ (11)
where $H(z)$ is the Hubble rate, $f_{s}$ is the frequency of the GW at the source, and $\mathcal{R}(z)$ is the merger rate of BBHs as a function of redshift. We have used [40]:
@@
×$\times p(m_{1},m_{2})\,,$ (12)
where we integrate the possible combinations of masses, $m_{1}$ and $m_{2}$, for a single binary. A significant amount of work is done to use individual events [51], or joint measurements of individual events and the GWB [40, 29, 52, 53] to constrain $p(m_{1},m_{2})$ and $\mathcal{R}(z)$. As a note, this curve is generated using popstock and intended to be an inflated BBH signal curve to emulate the magnitude of the O3 signal [35].
[…]
The merger rate as a function of redshift is often assumed to have a similar functional form to the star formation rate; the canonical argument being that the abundance of black hole binaries, and their subsequent mergers, should track the star formation rate with a time delay between formation and merger, subject to constraints due to the metallicity in the local environment which affects the birth rate of black holes [92, 93, 94, 95]. A common choice for this functional form is a smooth BPL similar to the Madau-Dickinson star formation rate, given by [40]
@@
where $R_{0}$ is the local merger rate, $z_{p}$ is the redshift at which the merger rate peaks, $\alpha$ the spectral index for $z\lesssim z_{p}$, $\beta$ is the negative of the spectral index for $z\gtrsim z_{p}$, and $\mathcal{C}$ is a normalization factor to ensure $\mathcal{R}(0)=R_{0}$. For our BBH-like injection, we assume $R_{0}=31.88\,{\rm Gpc}^{-3}{\rm yr}^{-1}$, $\alpha=1.9$, $\beta=3.4$, and $z_{p}=2.4$ [40].
The argument in favor of using a Madau-Dickinson-like merger rate broadly assumes that the bulk of mergers are from binaries formed in the field. This is in contrast to models like those formed dynamically in dense stellar clusters [96, 97], or other hypotheses such as lensed high redshift stellar-mass BBHs of $\mathcal{M}_{c}\lesssim 15M_{\odot}$ where $\mathcal{M}_{c}$ is the chirp mass of the binary [98].
@@
As stated previously, since the CE sensitivity injections are in the strong signal regime, Eq. 4 cannot be calculated by simply taking the PSD of each detector, as is often done currently. For this work, we assume that we have an independent measurement of $\sigma^{2}(f)$ (e.g. from a null stream like one might construct from a triangular detector like ET or LISA). However, a future extension of this work might be to fit $\sigma^{2}(f)$ (or the detector PSDs) with our flexible model simultaneously with the background, alleviating the need for an independent measurement.
@@
For A+ sensitivity, as shown in Fig. 2, each posterior has an average of 1 knot until 20 years of observation time, when the average increases to 2 knots. Signal detection is confirmed in less than 20 years with one knot, with a PL trend recovered after 20 years with 2 knots. Even with 1 knot for shorter observation times, the posteriors still reflect the injection’s PL behavior.
Since the number of knots in each posterior sample reflects whether a signal is present within a given sensitivity regime, we can use this information to construct a detection statistic. For example, the right hand side of Figure 2 shows that for $T_{\mathrm{obs}}>1\,\mathrm{yr}$ there are few if any posterior samples with $n=0$, which could indicate a potential detection. However, simply computing an odds ratio, the number of posterior samples with $n>0$ knots divided by those with $n=0$ knots, is not necessarily sufficient on its own. The prior odds built into the odds ratio inherently favor the presence of a signal over noise. A potentially more robust detection statistic might undo the inherent prior odds by dividing by the maximum number of knots $N_{\rm max}$, which produces the Bayes factor. However, the extra prior volume of including more available knots (higher $N_{\mathrm{max}}$) can result in a Bayes factor that inherently favors noise compared to signal. We discuss these trade-offs in more detail, along with the odds ratio and Bayes factor computations in the context of this transdimensional spline model, in Appendix B.
The convergence of the number of knots in the posteriors is reflected by the 95% confidence envelopes narrowing around the injection as we increase the observation time. For example, the envelopes for 1 and 5 years (green, orange) of observation time recover the injection for $f\lesssim 100$ Hz. Although we can fit the PL behavior of the injection for $f\lesssim 100$ Hz for a wide range of observation times, there is poor recovery at higher frequencies, where A+ detectors are less sensitive to the background signal.
@@
When comparing knot numbers for posteriors, the knots for CE sensitivity posteriors are better constrained per observation time than for A+ sensitivity. For CE sensitivity, on average $4-13$ knots are required to recover the injection, while A+ sensitivity requires anywhere from $1-15$ knots. Furthermore, as the observation time for CE sensitivity increases to 10 years, the range of knots decreases to $6-11$ knots, while the number of knots required for recovering injection with A+ sensitivity with 10 years of observation time is still $1-15$ knots. Better recovery at higher frequencies, implicated by fewer knots for injections with CE sensitivity, correspond to better sensitivity to late-inspiral corrections and the $\Omega_{\rm GW}(f)$ turnover. We conclude that whether we use A+ or CE sensitivity affects the convergence of the number of knots in each posterior, in addition to the accuracy of the recovery envelopes.
@@
Figure 3: The same injection and recovery as Fig. 2 but with a CE sensitivity curve. On the left, we show the 95% confidence recovery envelopes for the posteriors of the injected signal (solid black line). On the right, we show the number of knots in each posterior draw for each observation time as a heatmap. We are able to recover the injection well up to $f\sim 100$ Hz for all observing times with the number of knots converging as observation time increases.

### v2

@@ IV Flexible, Transdimensional Models
We propose an interpolation model to recover a GWB signal by placing a discrete number of points in an $n$-dimensional space and interpolating between these points. This requires a choice of interpolation scheme, which we describe in Sec. IV.1. The general interpolation model naturally lends itself to using a RJMCMC algorithm [greenReversibleJumpMarkov1995] that allows the number of knots and thus the number of parameters in our model to vary. So, we must define how these interpolating points, or knots, move for each proposal in the RJMCMC. These knot movements are our proposals and are defined in Sec. IV.2. This is similar in spirit to the BayesLine and BayesWave algorithms [Littenberg:2014oda, Cornish:2014kda, Cornish:2020dwh].
Similar interpolation schemes have been used to flexibly fit unknown spectral shapes in the past. For example, in [renziniImprovedLimitsStochastic2019, Xiao:2022uvq] the authors estimate maps of the GWB from LVK data independently in separate frequency bins, with a flexible binning scheme in [Xiao:2022uvq] based on the sensitivity to the GWB spectrum. A similar approach was presented in [capriniReconstructingSpectralShape2019a], where the authors employ frequency binning across the LISA spectrum using bins chosen based on the expected sensitivity curve. In [ghalebBayesianReconstructionPrimordial2025] the authors use a similar interpolation scheme to what we present below, with a variable number of knots. They choose the number of knots by repeatedly calculating the Bayesian evidence, while in this work we choose the knots dynamically with the reversible jump algorithm.
@@ IV.1 Interpolation models
Instead of using a specific functional form for the GWB, such as a PL, we instead estimate the GWB at a set of chosen frequencies, which we call “knots,” and interpolate from those knots to the frequency bins at which measurements are taken. The parameters of the model are a set of $n$ amplitudes $\mathbf{a}_{n}$, located at a set of knots $\mathbf{x}_{n}$, in frequency space. We use $x$ here to denote the location of a knot, to distinguish them from the frequency bins at which $\hat{C}(f)$ is calculated.
@@
with $f\in[x_{j},x_{j+1}]$. We refer to this model as the “linear interpolation” model or, if we perform the interpolation in log-space (i.e. $x\rightarrow\ln x$ and $a\rightarrow\ln a$ etc.) the “piecewise power law” (PPL) model because linear interpolation in log-space corresponds to constructing the model from a set of PPLs. By working in log-space we can ensure that the GWB is positive (a defining characteristic of $\Omega_{\textrm{GW}}$), and that in the case of $n=1$ or $n=2$ knots the model simplifies to a constant or PL model. Moreover, for $n=2$, placing a log-uniform prior on $\mathbf{a}$ means that when we recover a PL, the prior we use is naturally the preferred prior advocated for in [Callister:2017ocg] and used in [Callister:2017ocg, KAGRA:2021kbb].
@@
where the $c_{n,j}$’s are calculated using standard algorithms that solve for the $c_{n,j}$ subject to constraints that impose continuity and smoothness [deboor].
An additional interpolation method that we employ is the “Akima spline” model. Akima splines are nonsmoothing splines that avoid the rapid oscillations between knots that are common in other spline formulations when the data are uninformative or the second derivative of the underlying function varies rapidly. An Akima spline model takes the same form as the cubic spline, but the coefficients $c_{n,j}$ are informed by derivatives of neighboring knots instead of an optimization [WANG2014122]. The result of an Akima spline interpolation is a smoother interpolation that is less sensitive to outliers in the data.
@@ IV.2 Reversible jump Markov chain Monte Carlo
We start with an underlying grid of $N$ knots that are available to interpolate between. At any given time, $n$ of them will be used in the interpolation, or “activated,” and NnN-n knots will be disregarded in the interpolation, or “deactivated.” The locations of the available underlying knots are labeled $x_{\alpha}$, where $\alpha=0,\ldots,N-1$. So, the initial location of $x_{\alpha}$ is given by
@@
The parameters of our model are the list of amplitudes $\mathbf{a}_{n}$ corresponding to the amplitude of the model at the knot location $\mathbf{x}_{n}$ of each of the activated knots. We also allow the knot locations to vary but maintain ordering on the underlying grid. That is, $x_{j}$ can vary within the range $f_{\textrm{min}}+2j\Delta x<x_{j}<f_{\textrm{min}}+2(j+1)\Delta x$, where $\Delta x=(f_{\textrm{max}}-f_{\textrm{min}})/(2N)$. So at any given time we have $2n$ parameters in our model. The prior on $\mathbf{x}_{n}$ and $\mathbf{a}_{n}$ throughout this paper is uniform within their allowed regions, and disjoint between knots  [^2] This simply implies two knots cannot be superposed.:
$p(\mathbf{a}_{n},\mathbf{x}_{n})=(2\Delta x)^{-n}(a_{\textrm{max}}-a_{\textrm{min}})^{-n}.$ (10)
We discuss the effect of these priors and how they change as the number of knots changes in Appendix  A.
@@
    Birth proposal: We propose to “turn on” one knot that is currently deactivated, choosing from the list of deactivated knots at random. This increases the number of parameters in the model and the overall prior volume. The location and amplitude of the proposed knot is either drawn from the prior or the proposed $x_{j}^{new}$ is drawn from the prior, while $a_{j}$ is drawn from a Gaussian with a mean given by $\mathcal{G}(x_{j}^{new})$. In this case, the location and amplitudes of the knots for the interpolation, $x_{j}$ and $c_{n,j}$ in Eq. (6) or Eq. (8), are from the current sample. The standard deviation of the Gaussian is selected by the user when the sampler is instantiated (the default is $\sigma=1$).
@@
    Death proposal: We propose to “turn off” a knot that is currently active. This decreases the number of parameters in the model and the overall prior volume.
@@
    Shift knot horizontally: We choose an active knot at random and propose to shift its location $x_{j}$ within its allowable region.
@@
Figure  1 provides a schematic of the RJMCMC, illustrating a death proposal and visualizing the priors associated with each interpolating knot. We discuss the details of the algorithm in Appendix  A, including the basics of RJMCMC, and more details about the implementation of the jump proposals.
Figure 1: A visualization of the RJMCMC implemented in this work. In both panels, we show the injected signal (black dashed line) and the injected signal plus Gaussian white noise (blue points). The error bars on the blue points correspond to one standard deviation of the Gaussian white noise. In the top panel, the gray boxes correspond to the uniform priors that the interpolating knots (yellow stars) are allows to move within [see Eq. (10)]. The knots are interpolating using cubic splines (red lines). The bottom panel shows the death proposal (see Sec. IV.2) where we have turned off one knot (gray star) and interpolated the remaining knots to fit the signal. This schematic highlights the transdimensional nature of the RJMCMC that we implement to fit injected GW signals with various detector sensitivities.
@@ V Detecting and Characterizing $\Omega_{\rm GW}(f)$
We choose three example injections to demonstrate our transdimensional spline interpolation model. These injections include a BBH-induced signal (Sec. V.1), a subtraction signal (Sec. V.2), and a signal based on a first-order phase transition model (Sec. V.3). For each example, after motivating the astrophysical relevance of each injection, we present the results of our injection and recovery. We perform the injection and recovery with both LIGO A+ and projected Cosmic Explorer (CE) sensitivity curves. For the CE sensitivity curve, we assume 40 km detector arms and the same location and orientation as the current LIGO interferometers at Hanford, WA and Livingston, LA [Evans:2021gyd, Evans:2023euw].
Before performing the injection study, we present the cumulative signal-to-noise ratio (SNR) curves for the three different injections considered in this work in Fig. 2, where the SNR for stochastic signals is defined as the optimal filter SNR of [Allen:1999]. Vertical lines mark the frequency values that correspond to where 99% of the SNR is accumulated. As may be observed, in all cases frequencies above 40 Hz contribute less than 1% of the SNR.
For each injection, we simulate $\hat{C}(f)$ directly, by drawing $\hat{C}(f)\sim\mathcal{N}[\Omega_{\mathcal{M}}(f),\sigma(f)]$, where $\Omega_{\mathcal{M}}(f)$ is the injected model, and $\sigma(f)$ is the detector noise calculated using the specified noise curves and observation times. For each recovery, we perform 500 000500\,000 Markov chain Monte Carlo steps, and discard the first 300 000300\,000 as burn-in. We perform all interpolation in log-space; i.e. we take $x\rightarrow\ln x$, $a\rightarrow\ln a$, etc., when implementing the procedures outlined in Secs. IV.1 and IV.2.
Figure 2: Cumulative signal-to-noise ratio (SNR) curves for the three different injections considered in this work: BBH background (purple), subtraction background (red), and first-order phase transition (FOPT) signal (brown). For reference, we include the scaling of the cumulative SNR on the observation time, for the BBH signal case: 1 month (dotted line), 1 year (full line), and 2 years (dashed line). This dependence scales identically for the different signals. The SNR for stochastic signals employed here is defined as the optimal filter SNR in [Allen:1999]. Vertical lines mark the frequency values that correspond to where 99% of the SNR is accumulated.
@@ V.1 Full BBH injection
We use a BBH $\Omega_{\rm GW}(f)$ as our first example of injection and recover it using this transdimensional spline interpolation approach. The contribution to the GWB from all BBHs can be written as
$\Omega_{\textrm{GW}}(f)=\frac{f}{\rho_{c}}\int_{0}^{z_{\textrm{max}}}\textrm{d}z\;\frac{\mathcal{R}(z)\left<\frac{\textrm{d}E}{\textrm{d}f_{s}}\left.\right|_{f(1+z)}\right>}{(1+z)H(z)}\,$, (11)
where $H(z)$ is the Hubble rate, $f_{s}$ is the frequency of the GW at the source, and $\mathcal{R}(z)$ is the merger rate of BBHs as a function of redshift. We have used [Callister:2020arv]
$\bigg\langle\frac{dE}{df_{s}}|_{f(1+z)}\bigg\rangle=\int dm_{1}dm_{2}\frac{dE}{df_{s}}$ $\big(m_{1},m_{2};f(1+z)\big)$
$\times p(m_{1},m_{2})\,$, (12)
where we integrate the possible combinations of masses $m_{1}$ and $m_{2}$ for a single binary. A significant amount of work is done to use individual events [KAGRA:2021duu], or joint measurements of individual events and the GWB [Callister:2020arv, KAGRA:2021kbb, Turbang:2023tjk, Lalleman:2025xcs] to constrain $p(m_{1},m_{2})$ and $\mathcal{R}(z)$, where $p(m_{1},m_{2})$ is the probability distribution of the source masses. As a note, this curve is generated using POPSTOCK and intended to be an inflated BBH signal curve to emulate the magnitude of the O3 signal [Renzini:2024pxt].
@@
The merger rate as a function of redshift is often assumed to have a similar functional form to the star formation rate (SFR); the canonical argument being that the abundance of black hole binaries, and their subsequent mergers, should track the star formation rate with a time delay between formation and merger, subject to constraints due to the metallicity in the local environment which affects the birth rate of black holes [Regimbau:2011rp, Rosado:2011kv, LIGOScientific:2016fpe, Callister:2016ewt]. A common choice for this functional form is a smooth BPL similar to the Madau-Dickinson star formation rate, given by [Callister:2020arv]
@@
where $R_{0}$ is the local merger rate, $z_{p}$ is the redshift at which the merger rate peaks, $\alpha$ is the spectral index for $z\lesssim z_{p}$, $\beta$ is the negative of the spectral index for $z\gtrsim z_{p}$, and $\mathcal{C}$ is a normalization factor to ensure $\mathcal{R}(0)=R_{0}$. For our BBH-like injection, we assume $R_{0}=31.88\,{\rm Gpc}^{-3}{\rm yr}^{-1}$, $\alpha=1.9$, $\beta=3.4$, and $z_{p}=2.4$ [Callister:2020arv].
The argument in favor of using a Madau-Dickinson-like merger rate broadly assumes that the bulk of mergers are from binaries formed in the field. This is in contrast to models like those formed dynamically in dense stellar clusters [Mapelli_2021, Mandel_2022], or other hypotheses such as lensed high-redshift stellar-mass BBHs of $\mathcal{M}_{c}\lesssim 15\,M_{\odot}$ where $\mathcal{M}_{c}$ is the chirp mass of the binary [Dominik_2013].
@@
As stated previously, since the CE sensitivity injections are in the strong signal regime, Eq. (4) cannot be calculated by simply taking the PSD of each detector, as is often done currently. For this work, we assume that we have an independent measurement of $\sigma^{2}(f)$ (e.g. from a null stream like one might construct from a triangular detector like ET or LISA). However, a future extension of this work might be to fit $\sigma^{2}(f)$ (or the detector PSDs) with our flexible model simultaneously with the background, alleviating the need for an independent measurement.
@@
For A+ sensitivity, as shown in Fig. 3, each posterior has an average of one knot until 20 years of observation time, when the average increases to two knots. Signal detection is confirmed in less than 20 years with one knot, with a PL trend recovered after 20 years with two knots. Even with one knot for shorter observation times, the posteriors still reflect the injection’s PL behavior.
Since the number of knots in each posterior sample reflects whether a signal is present within a given sensitivity regime, we can use this information to construct a detection statistic. For example, the right-hand side of Fig. 3 shows that for $T_{\textrm{obs}}>1\,\textrm{yr}$ there are few if any posterior samples with $n=0$, which could indicate a potential detection. However, simply computing an odds ratio, the number of posterior samples with $n>0$ knots divided by those with $n=0$ knots, is not necessarily sufficient on its own. The prior odds built into the odds ratio inherently favor the presence of a signal over noise. A potentially more robust detection statistic might undo the inherent prior odds by dividing by the maximum number of knots $N_{\rm max}$, which produces the Bayes factor. However, the extra prior volume of including more available knots (higher $N_{\textrm{max}}$) can result in a Bayes factor that inherently favors noise compared to signal. We discuss these trade-offs in more detail, along with the odds ratio and Bayes factor computations in the context of this transdimensional spline model, in Appendix B.
The convergence of the number of knots in the posteriors is reflected by the 95% confidence envelopes narrowing around the injection as we increase the observation time. For example, the envelopes for 1 and 5 years (green and orange, respectively) of observation time recover the injection for $f\lesssim 100$ Hz. Although we can fit the PL behavior of the injection for $f\lesssim 100$ Hz for a wide range of observation times, there is poor recovery at higher frequencies, where A+ detectors are less sensitive to the background signal.
@@
In Figs. 3 and 4, we see that in regions of the spectrum where the data are uninformative, the model is extrapolated from the parts of the spectrum where is more constrained. In Fig. 3, this results in a mismatch around the merger-induced peak in the spectrum for the 10 yr case. For Fig. 4, for 1 month we are unable to capture the late-inspiral-induced “dip” in the spectrum around $50$ Hz because 99% of the SNR has been accumulated by $\sim 35$ Hz (cf. Fig. 2). Instead, the power law constrained at low frequencies is extrapolated to higher frequencies.
When comparing knot numbers for posteriors, the knots for CE sensitivity posteriors are better constrained per observation time than for A+ sensitivity. For CE sensitivity, on average 44-$13$ knots are required to recover the injection, while A+ sensitivity requires anywhere from 11-$15$ knots. Furthermore, as the observation time for CE sensitivity increases to 10 years, the range of knots decreases to 66-$11$ knots, while the number of knots required for recovering injection with A+ sensitivity with 10 years of observation time is still 11-$15$ knots. Better recovery at higher frequencies, implicated by fewer knots for injections with CE sensitivity, corresponds to better sensitivity to late-inspiral corrections and the $\Omega_{\rm GW}(f)$ turnover. We conclude that whether we use A+ or CE sensitivity affects the convergence of the number of knots in each posterior, in addition to the accuracy of the recovery envelopes.
@@
Figure 4: The same injection and recovery as Fig. 3 but with a CE sensitivity curve. On the top left, we show the 95% confidence recovery envelopes for the posteriors of the injected signal (solid black line). The bottom left panel shows the residuals, $\Omega_{\rm GW,res}$, which are obtained by subtracting the 9595% confidence intervals in the top left panel from the injected signal. On the right, we show the number of knots in each posterior draw for each observation time as a heat map. We are able to recover the injection well up to $f\sim 100$ Hz for all observing times with the number of knots converging as observation time increases.